OpenStudy (anonymous):

in a chemical process, the flow rate of a soulution is 145 liters per second. Epress this rate in liters per minute

OpenStudy (anonymous):

145 litres => 1/60 minutes X litres => 1 minute solve for X :)

OpenStudy (anonymous):

i got that but im lost after that

OpenStudy (anonymous):

cross multiply the equations . 145 ( 1) = X ( 1/60)

OpenStudy (anonymous):

=> that X = (60)(1)(140) = … litres/min

OpenStudy (anonymous):

145*60=8700 liter pm

OpenStudy (anonymous):

that ^